Integration by parts

1. Integral of 5x sin4x dx

2. Integral of x lnx dx
1
Expert's answer
2019-12-10T10:57:26-0500

Calculus

We need to find the Integrations using by parts


Solution:


"\int f(x) \space g(x) dx = f(x) \int g(x) dx -\int f' (x) (\int g(x) dx ) dx"


1)


"\int 5x \space sin \space 4x \space dx = 5 \int x \space sin \space 4x \space dx"


Here, f(x) = x and g(x) = sin 4x


Then "f ' (x) = 1"


"\int 5x \space sin \space 4x \space dx =5 \Bigg( x \int sin 4x - \int 1. (\int sin 4x dx ) dx \Bigg)"



"= 5 \Bigg( x (\frac {-cos 4x }{4}) - \int (\frac {-cos 4x }{4}) dx \Bigg) + c"



"= 5 \Bigg( \frac {- x cos 4x}{4 } + \frac {sin 4x}{16} \Bigg) + c"

2)


"\int x \space ln x \space dx = ln x \int x \space dx - \int d( ln x)( \int x dx ) dx"



"= ln x \times \frac {x^2 }{2} - \int \frac {1}{x} \times \frac {x^2}{2} dx + c"

"= \frac {x^2}{2} \times ln x - \frac {1}{2} \int x dx + c"


"= \frac {x^2}{2} \times ln x - \frac {1}{2} \times \frac {x^2} {2} + c"

"= = \frac {x^2}{2} \times ln x - \frac {x^2} {4} + c"